# Dark monopoles

### v2

@@ Abstract
Monopoles have been a subject of much theoretical and experimental research since they were proposed to symmetrize Maxwells equations. However, no experimental signature of their existence has been detected. Many mechanisms have been proposed to explain this lack of success. In here we generalize QED to a two photon theory where the magnetic photon and the monopole belong to dark matter. A naive mixing interaction produces duality and generates experimental consequences for indirect monopole detection via the magnetic charge of the electron.
@@ 1 Introduction
Dirac’s beautiful interpretation of the quantum mechanical behavior of the magnetic monopole led to a technical difficulty associated with the famous Dirac string [1, 2]. The formalism of Yang and Wu [3, 4] using fiber bundles and the two photon theory description by Cabibbo and Ferrari [5] simplified the description. In our case we start by defining a dual Maxwell theory using a two photon approach [5, 6, 7, 8], where we consider the magnetic photon and the monopole as dark matter [9, 10]. We next assume that dark and ordinary sectors interact through a portal in a manner that, though feeble, is experimentally accessible. In our portal the interaction takes place because of the kinetic mixing between the dark and the visible photons [11, 12]. A description which has similar motivation to ours is SO(2) dual electrodynamics of electrically and magnetically charged particles in interaction [13]. We study for simplicity the theory for electron and monopole. In the most general scenario both the electron and the monopole become dyons acquiring electric and magnetic charge. We propose several experiments to maximize the effect of these charges and describe their observational consequences.
@@ 2 Dark photons
where $\mathcal{F}_{\mu\nu}=\partial_{\mu}\mathcal{A}_{\nu}-\partial_{\nu}\mathcal{A}_{\mu}$ and $\mathcal{G}_{\mu\nu}=\partial_{\mu}\mathcal{C}_{\nu}-\partial_{\nu}\mathcal{C}_{\mu}$. We depart from Cabibbo and Ferrari in that we will ascribe the monopoles to the dark matter. Thus, $\mathcal{A}_{\mu}$ is the electric photon that couples only to particles with electric charge and $\mathcal{C}_{\mu}$ the magnetic photon that couples only to particles with magnetic charge. The interaction lagrangian with matter is
@@
where $J^{e}_{\mu}$ is the electric current and $J^{m}_{\mu}$ the magnetic current. For simplicity we will consider visible matter to be made just of electrons and dark matter just of monopoles. We assume both photons to be massless. In order to generate a portal we assume a kinetic mixing between the photons
@@
This kinetic term can be diagonalized by rotating the photons [11, 12]
(${\mathcal{C}^{\mu}\choose\mathcal{A}^{\mu}}=\left(\begin{array}[]{cc}{\displaystyle\frac{1}{\sqrt{1-\varepsilon^{2}}}}&0\\ {\displaystyle-\frac{\varepsilon}{\sqrt{1-\varepsilon^{2}}}}&1\end{array}\right)\left(\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right){C^{\mu}\choose A^{\mu}}\,,$ (8)
where now we can identify $A^{\mu}$ with the visible photon and $C^{\mu}$ with the dark photon. After this rotation the photon lagrangian becomes diagonal
@@
This equation is very illuminating. It shows that in the rotated theory the electron becomes a dyon with electric and magnetic charge
@@ 3 Experimental consequences
which has been plotted in Fig. 1. The smallness of the magnetic charge effects in the electron imply that the physical region must be $\theta\sim 0$
@@
Let us discuss two possibilities satisfying the experimental restrictions:
@@ 6 Production of monopoles
Figure 5: The structure of the total cross sections for $e^{+}e^{-}$ annihilation into monopole-antimonople: for $S=0$(dashed) and for $S=\nicefrac{{1}}{{2}}$ (solid). We use here $(2\theta-\varepsilon)^{2}=1$ and the cross section in units of $m^{-2}$.
@@
We next calculate the maximun cross section for monopole masses $10$ GeV $<m<10000$ GeV and study the corresponding dependence with $s$. We divide this result by the cross section $e^{+}e^{-}\rightarrow\mu^{+}\mu^{-}$, and by that of $e^{+}e^{-}\rightarrow hadrons$. The agreement between theory and experiment for $e^{+}e^{-}\rightarrow\mu^{+}\mu^{-}$ is excellent ($<1$% level) across a wide range of energies, validating QED and the Standard model [22, 23, 24, 25]. For $e^{+}e^{-}\rightarrow hadrons$ the agreement is excellent in the high energy region ($<1$% level); the agrement is good in the intermediate-energy range ($<5$% level) and poorer agreement is reached in the low energy region limited by non-perturbative QCD and hadronic resonance modeling [24, 25]. In our case we are interested in the high energy region where the experimental precission to produce monopoles has to be at the maximum at the $1$% level. By using this limit we find upper bounds for the coupling constant appearing in Eq.(33) and Eq.(34), $(2\theta-\varepsilon)^{2}$, which we show in Fig.6. The extremely good agreement for these cross sections in the high energy region between experiment and theory restricts greatly the region of the couplings. All the cross sections entering the analysis have been calculated to lowest order, that is the reason behind the absence of $s$ dependence in the ratios.
@@
Figure 6: The ratio of the monopole production cross section to $\mu$ pairs (dotted) and to hadrons (dot-dashed) divided by 100 for $S=0$ monopoles , and the same for muon pairs (solid) and to hadrons (dotted) divided by 100 for $S=\nicefrac{{1}}{{2}}$ monopoles.
@@
Next we plot in the left hand side of Fig. 7 the maximum monopole production cross section as a function of center of mass energy for 0.1 and 0.001 times the limiting coupling constants, both for the $S=0$ case (dashed) and for the $S=\nicefrac{{1}}{{2}}$ (solid). Recall that the maximum cross section occurs for $s=10m^{2}$ in the scalar case and $s=6m^{2}$ in the fermion case (see Fig. 5), thus each value of the center-of-mass energy has an associated monopole mass. In these ranges of couplings the effect is small but measurable, similar to that obtained by calculations in other schemes [26] as we show in the right hand side of Fig. 7, where we compare the present results with those of the calculation of ref. [26] for the same energy range. It is immediately apparent from our results that the mass of the monopole plays a crucial role. The smaller the mass the larger the detection cross section will be. However, there is a big difference between previous calculations and the present one, namely in the previous analysis the monopole couples strongly to real photons and therefore the monopole radiates real photons or monopole-antimonopole annihilate into photons [26, 27]. In the present model the monopole couples only weekly to real photons and therefore both processes are further suppressed. This is the reason for not having detected the monopole yet, namely its properties as a dark matter particle.
@@ 7 Conclusions
We have described an electromagnetic theory for electrons and monopoles with two photons. We assume the magnetic sector of the theory to be associated with dark matter. We propose a naive coupling between the two photons and diagonalize in order to separate the real photon from the dark photon. In this process the electron acquires a magnetic charge and the monopole an electric charge. We have analyzed two limiting scenarios characterized by the value of the parameter $\theta$. In the case $\theta=0$ the monopole remains a dark monopole decoupled from the real world, while the electron acquires a very small magnetic charge. This magnetic charge inspires experimental ways of investigating the ideas behind this approach. We have performed two analysis to see how the magnetic charge of the electron could be determined. The first experimental proposal is a generalization of a Biot-Savart experiment, while the second generalizes the Aharonov-Bohm experiment. These two proposals benefit from the peculiarity of the electron in this model, which becomes a dyon with a very tiny magnetic charge.
In the other scenario $\sin\theta=\varepsilon$ the monopole acquires a very small electric charge, and therefore production in accelerators is out of the question. Some implications of our results in astrophysics might be possible. If we relax the latter coupling accepting a smaller value of the angle, we have a model with both electric and magnetic charges for electron and monopoles. In this case the first two experimentl proposal are also feasible and moreover, we can produce monopoles. The production cross section will be proportional to $e^{2}g^{2}$ times a small portal coupling, either $\theta$ or $\theta-\varepsilon$, therefore the model predicts a small production cross section. This might be the reason for not having been able to detect the monopole. We have though studied the possible upper limits of the mixing couplings by comparing the production by electron-positron annihilation of a monopole-antimonopole pair with a $\mu^{+}-\mu^{-}$ pair or hadrons below the limit of precission of the latter. We see that these upper limits although small are not completely unrealistic. However, the fact that monopoles couple weakly to real photons makes their detection further suppressed.
@@ Acknowledgments
This work was inspired by conversations with Huner Fanchiotti and Carlos García Canal. VV acknowledges illuminating email exchanges with D. Singleton regarding the two photon model, and Jan Govaerts regarding the SO(2) model. VV would like to thank Lukas Theussl for help with JaxoDraw [28]. Support by Grant CNS2022-135688 Consolidación Investigadora is acknowledged.

### v3

@@ Abstract
Monopoles have been a subject of much theoretical and experimental research since they were proposed to symmetrize Maxwell’s equations. However, no experimental signature of their existence has been detected. Many mechanisms have been proposed to explain this lack of success. In here we generalize QED to a two photon theory where the extraordinary photon and the monopole belong to the dark sector, and the ordinary photon and the electron to the non-dark sector. A mixing interaction transforms both electron and monopole into dyons generating experimental consequences for indirect and direct monopole detection which we analyze.
@@ 1 Introduction
Dirac’s beautiful interpretation of the quantum mechanical behavior of the magnetic monopole led to a technical difficulty associated with the famous Dirac string [1, 2]. The formalism of Yang and Wu [3, 4] using fiber bundles and the two photon theory of Cabibbo and Ferrari [5] simplified the description avoiding strings. In here we present a model of two photons [5, 6, 7, 8], containing an ordinary photon interacting only with electric currents and an extraordinary photon interacting only with magnetic currents, the latter in the dark sector [9, 10]. We next assume that the dark and ordinary sectors interact through a portal in a manner that, though feeble, is experimentally accessible. In our portal the interaction takes place because of the kinetic mixing between the dark and the ordinary photons [11, 12]. Our model differs from that of ref. [11] because we work in QED, not in the standard model, to which we add an additional $U(1)$ dark photon à la Cabibbo and Ferrari. In the most general scenario of this model, both electrons and monopoles become dyons, acquiring electric and magnetic charges. Based of this result we propose several experiments to maximize the effect of the new charges, and their observational consequences. A description which has similar motivation to ours is SO(2) dual electrodynamics of electrically and magnetically charged particles in interaction [13].
@@ 2 Dark photons
where $\mathcal{F}_{\mu\nu}=\partial_{\mu}\mathcal{A}_{\nu}-\partial_{\nu}\mathcal{A}_{\mu}$ and $\mathcal{G}_{\mu\nu}=\partial_{\mu}\mathcal{C}_{\nu}-\partial_{\nu}\mathcal{C}_{\mu}$. We depart from Cabibbo and Ferrari in that we will ascribe the monopoles to dark matter. Thus, $\mathcal{A}_{\mu}$ is the ordinary photon that couples only to particles with electric charge and $\mathcal{C}_{\mu}$ the extraordinary photon that couples only to particles with magnetic charge. The interaction lagrangian with matter is
@@
where $J^{e}_{\mu}$ is the electric current and $J^{m}_{\mu}$ the magnetic current. For simplicity we will consider ordinary matter to be made just of electrons and dark matter just of monopoles. We assume both photons to be massless. In order to generate a portal we assume a kinetic mixing between the photons
@@
This kinetic term can be diagonalized by transforming the photons [11, 12]
${\mathcal{C}^{\mu}\choose\mathcal{A}^{\mu}}=\left(\begin{array}[]{cc}{\displaystyle\frac{1}{\sqrt{1-\varepsilon^{2}}}}&0\\ {\displaystyle-\frac{\varepsilon}{\sqrt{1-\varepsilon^{2}}}}&1\end{array}\right)\left(\begin{array}[]{cc}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right){C^{\mu}\choose A^{\mu}}\,$, (8)
where now we identify $A^{\mu}$ with the visible photon and $C^{\mu}$ with the dark photon. We keep the nomenclature dark photon due to its origin despite the fact that it might couple weakly to ordinary particles as we explain next.
After this transformtion the photon lagrangian becomes diagonal
@@
This equation is very illuminating. It shows that in the transformed theory the electron becomes a dyon with electric and magnetic charge
@@
Thus the dark photon gets a coupling to the electron and the ordinary photon to the monopole thanks to the mixing term. The new couplings depend not only on the parameter $\varepsilon$, a coupling constant, but also on the parameter $\theta$. For $sin\theta=0$ the ordinary photon must couple to the electron as in ordinary QED. This fixes the normalization of the fields and therefore different values of $\theta$ lead to different charges, thus to different physical scenarios [11, 12]. The aim of the following sections is to discuss various interesting physical scenarios associated with various values for the charges.
@@ 3 Experimental consequences
which has been plotted in Fig. 1. The smallness of the magnetic charge effects in the electron imply that the physical region must be $\theta\sim 0$ .
@@
Let us discuss two possibilities satisfying the experimental restriction:
@@ 6 Production of monopoles
Figure 5: The structure of the total cross sections for $e^{+}e^{-}$ annihilation into monopole-antimonople: for S=0S=0(dashed) and for $S=\nicefrac{{1}}{{2}}$ (solid). We use here $(2\theta-\varepsilon)^{2}=1$ and the cross section in units of $m^{-2}$.
@@
We next calculate the maximun cross section for monopole masses $10$ GeV <m<10000<m<10000 GeV and study the corresponding dependence with $s$. We divide this result by the cross section $e^{+}e^{-}\rightarrow\mu^{+}\mu^{-}$, and by that of $e^{+}e^{-}\rightarrow hadrons$. The agreement between theory and experiment for $e^{+}e^{-}\rightarrow\mu^{+}\mu^{-}$ is excellent (<1<1% level) across a wide range of energies, validating QED and the Standard model [21, 22, 23, 24]. For $e^{+}e^{-}\rightarrow hadrons$ the agreement is excellent in the high energy region (<1<1% level); the agrement is good in the intermediate-energy range (<5<5% level) and poorer agreement is reached in the low energy region limited by non-perturbative QCD and hadronic resonance modeling [23, 24]. In our case we are interested in the high energy region where the experimental precission to produce monopoles has to be at the maximum at the 11% level. By using this limit we find upper bounds for the coupling constant appearing in Eq.(33) and Eq.(34), $(2\theta-\varepsilon)^{2}$, which we show in Fig.6. The extremely good agreement for these cross sections in the high energy region between experiment and theory restricts greatly the region of the couplings. All the cross sections entering the analysis have been calculated to lowest order, that is the reason behind the absence of $s$ dependence in the ratios.
@@
Figure 6: The ratio of the monopole production cross section to $\mu$ pairs (dotted) and to hadrons (dot-dashed) for $S=0$ monopoles , and the same for muon pairs (solid) and to hadrons (dotted) for $S=\nicefrac{{1}}{{2}}$ monopoles all ratios multiplied by $10^{2}$.
@@
Next we plot in the left hand side of Fig. 7 the maximum monopole production cross section as a function of center of mass energy for 0.1 and 0.001 times the limiting coupling constants, both for the $S=0$ case (dashed) and for the $S=\nicefrac{{1}}{{2}}$ (solid). Recall that the maximum cross section occurs for $s=10m^{2}$ in the scalar case and $s=6m^{2}$ in the fermion case (see Fig. 5), thus each value of the center-of-mass energy has an associated monopole mass. In these ranges of couplings the effect is small but measurable, similar to that obtained by calculations in other schemes, as we show in the right hand side of Fig. 7, where we compare the present results with those of the calculation of ref. [25] for the same energy range. It is immediately apparent from our results that the mass of the monopole plays a crucial role. The smaller the mass the larger the detection cross section will be.
There is a huge difference between the previous ordinary matter calculations and the present one, namely in previous analysis the monopole couples with electromagnetic strength to ordinary photons, and therefore the monopoles can be detected by their radiation to ordinary photons or the monopole-antimonopole pairs can be detected by their annihilation into ordinary photons [25, 26], this is not the case here, since the monopole and the monopole-antimonopole pairs couple weakly to ordinary photons. Thus our calculation favors direct detection of monopoles (antimonopoles) [27] over the detection via photon radiation.
@@ 7 Conclusions
We have described an electromagnetic theory for electrons and monopoles with two photons. We assume the magnetic sector of the theory to be associated with a dark sector. We propose a naive coupling between the two photons and diagonalize the interaction in order to separate the ordinary photon from the dark photon. In this process the electron acquires a magnetic charge and the monopole an electric charge. We have analyzed two limiting scenarios characterized by the value of the parameter $\theta$. In the case $\theta=0$ the monopole remains a dark monopole decoupled from the real world, while the electron acquires a very small magnetic charge. This magnetic charge inspires experimental ways of investigating the ideas behind this approach. We have performed two analysis to see how the magnetic charge of the electron could be determined. The first experimental proposal is a generalization of a Biot-Savart experiment, while the second generalizes the Aharonov-Bohm experiment. These two proposals benefit from the peculiarity of the electron in this model, which becomes a dyon with a very tiny magnetic charge.
In the other scenario $\sin\theta=\varepsilon$ the monopole acquires a very small electric charge, and therefore production in accelerators is out of the question. Some implications of our results in astrophysics might be possible. If we relax the latter coupling accepting a smaller value of the angle, we have a model with both electric and magnetic charges for electron and monopoles. In this case the two previous experimental proposals are also feasible, and moreover, we can produce monopoles. The production cross section is proportional to $e^{2}g^{2}$ times a small portal coupling, either $\theta$ or $\theta-\varepsilon$, therefore the model predicts a small production cross section. This might be the reason for not having been able to detect monopoles. We have studied the possible upper limits of the couplings by comparing the production by electron-positron annihilation of monopole-antimonopole pairs to $\mu^{+}\mu^{-}$ pairs (or hadrons) below the limit of precision of the latter. We see that these upper limits although small are not completely negligible. However, the fact that dark monopoles couple weakly to ordinary photons makes their detection via their decay into ordinary photons further suppressed. Direct detection of monopoles is favored over detection via their radiation.
@@ Acknowledgments
This work was inspired by conversations with Huner Fanchiotti and Carlos García Canal. VV acknowledges illuminating email exchanges with D. Singleton regarding the two photon model, Jan Govaerts regarding the SO(2) model, Marco Fabbrichesi regarding some aspects of ref. [12], and Arcadi Santamaria regarding field redefinitions. VV would like to thank Lukas Theussl for help with JaxoDraw [28]. Support by Grant CNS2022-135688 Consolidación Investigadora is acknowledged.
Data Availability Statement
This manuscript has no associated data since all the data generated appear in the figures.